# Deep Reinforcement Learning for Wireless Scheduling in Distributed Networked Control

### v4

@@ V-C Action Embedding for DDPG and TD3
To further handle the issue of large discrete action space, the discrete actions can be embedded into a continuous action space. This approach uses a vector of continuous values to represent the priority of allocating channels to plants. Specifically, each plant is assigned a continuous value. Plants are then allocated channels based on the ranking of these continuous values. For instance, the plant with the highest value is allocated to the first channel, the second highest to the second channel, and so on. This method significantly reduces the complexity of the discrete action space, making it feasible to handle larger $N$ and $M$. For example, the number of output neurons is $6$ when $N=6$ and $M=4$, while the original discrete action space is $1045$. The continuous action space allows for more efficient learning and optimization by leveraging the capabilities of DRL algorithms designed for continuous spaces. We will illustrate numerical results of using DRL for solving the scheduling problem in the following section.
@@ VI Numerical Results
We adopt the following hyper-parameters for the DRL algorithms. The input state dimension of DQN is set to $2N(v_{i}+1)$. We use three hidden layers with $300$, $200$, $100$ neurons, respectively. The output layer of DQN has $|\tilde{\mathcal{A}}|=\sum_{m=0}^{M}\mathsf{C}^{m}_{M}\mathsf{P}^{m}_{N}$ outputs (actions) for the problem with reduced action space (see Section V-B). The activation function in each hidden layer is the ReLU [49]. The activation function in the output layer of DDPG and TD3 is the Sigmoid [49]. The experience reply memory has a size of $K=100000$, and the size of mini-batch is $64$. So in each time step, $64$ data is sampled from the reply memory for training. The exploration parameter $\epsilon$ of DQN decreases from $1$ to $0.01$ at the rate of $0.999$ after each time step. The exploration method for DDPG and TD3 is based on Gaussian action noise with a standard deviation of $0.2$ and a mean value of 00. We adopt the ADAM optimizer to update the DNNs. Each training takes $E=500$ episodes, each including $T=500$ time steps. We test the trained policy for $100$ episodes and compare the empirical average costs $\frac{1}{T}\sum_{k=1}^{T}\sum_{i=1}^{N}c(\mathbf{s}_{i,k})$.
We first consider a $N$-plant-$M$-frequency WNCS with $N=5$ and $M=5$, i.e., $10$ links sharing $5$ frequencies with a reduced discrete action space of $1546$. Then, we will extend the network to a larger scale with $N=10$ and $M=10$, i.e., $20$ links sharing $10$ frequencies with a reduced discrete action space of $234662231$.
@@
In addition, Fig. 3 demonstrates that the well-trained DQN-based policy reduces the long-term average cost of greedy policy by $9.5$% from $826.2$ to $747.9$. The DDPG and TD3 algorithms based on the action embedding method achieve similar or slightly better performance compared to the DQN-based algorithm. This is attributed to the large discrete action space, which makes it difficult for the DQN-based algorithm to explore adequately during training and to get the optimal policy after training. The DDPG algorithm reduces the long-term average cost of the greedy policy by $14.5$% from $826.2$ to $706.0$. TD3 is an extension of DDPG, which is designed to reduce overestimation bias and variance in the learning process by introducing delayed policy updates, target policy smoothing, and twin critic networks. These features can lead to overly conservative behaviour in WNCS with high stochasticity, which potentially slow down the learning process or lead to underfitting where the system fails to adequately capture the beneficial actions amidst noise. Thus, DDPG, as a more aggressive approach than TD3.
@@
Fig. 4 showcases the comparison of three DRL algorithms and the three heuristic benchmark policies when $N=10$ and $M=10$. In this case, the DQN-based algorithm is unable to solve the MDP, as the reduced discrete action space is still too large. Results show that the DDPG and TD3 algorithms based on the action embedding method effectively address the challenge of the rapidly growing action space, making it more suitable for larger-scale WNCS problems. The DDPG algorithm reduces the long-term average cost of the greedy policy by $9.9$% from $542.7$ to $602.0$.
@@ VII Conclusions
We have investigated the transmission scheduling problem of the fully distributed WNCS. A sufficient stability condition of the WNCS in terms of both the control and communication system parameters has been derived. We have proposed advanced DRL algorithms to solve the scheduling problem, which performs much better than benchmark policies. For future work, we will develop a distributed DRL approach for enhancing the scalability of the scheduling algorithm. Furthermore, we will consider co-design problems of the estimator, the controller, and the scheduler for distributed WNCSs with time-varying channel conditions.

### v5

@@ V-C Action Embedding for DDPG and TD3
To further handle the issue of large discrete action space, the discrete actions can be embedded into a continuous action space. This approach uses a vector of continuous values to represent the priority of allocating channels to plants. Specifically, each plant is assigned a continuous value. Plants are then allocated channels based on the ranking of these continuous values. For instance, the plant with the highest value is allocated to the first channel, the second highest to the second channel, and so on. This method significantly reduces the complexity of the discrete action space, making it feasible to handle larger $N$ and $M$. For example, the number of output neurons is $6$ when $N=6$ and $M=4$, while the original discrete action space is $1045$. The continuous action space allows for more efficient learning and optimization by leveraging the capabilities of DRL algorithms designed for continuous spaces. We will illustrate the numerical results of using DRL for solving the scheduling problem in the following section.
@@ VI Numerical Results
VI-A Experiment Setup
We adopt the following hyper-parameters for the DRL algorithms. The input state dimension of DQN is set to $2N(v_{i}+1)$. We use three hidden layers with $300$, $200$, $100$ neurons, respectively. The output layer of DQN has $|\tilde{\mathcal{A}}|=\sum_{m=0}^{M}\mathsf{C}^{m}_{M}\mathsf{P}^{m}_{N}$ outputs (actions) for the problem with reduced action space (see Section V-B). The activation function in each hidden layer is the ReLU [57]. The activation function in the output layer of DDPG and TD3 is the Sigmoid [57]. The experience replay memory has a size of $K=100000$, and the mini-batch size is $64$. So in each time step, $64$ data is sampled from the replay memory for training. The exploration parameter $\epsilon$ of DQN decreases from $1$ to $0.01$ at the rate of $0.999$ after each time step. The exploration method for DDPG and TD3 is based on Gaussian action noise with a standard deviation of $0.2$ and a mean value of 00. We adopt the ADAM optimizer to update the DNNs. Each training takes $E=500$ episodes, each including $T=500$ time steps. We test the trained policy for $100$ episodes and compare the empirical average costs $\frac{1}{T}\sum_{k=1}^{T}\sum_{i=1}^{N}c(\mathbf{s}_{i,k})$.
VI-B Performance Comparison
@@
In addition, Fig. 3 demonstrates that the well-trained DQN-based policy reduces the long-term average cost of greedy policy by $9.5$% from $826.2$ to $747.9$. The DDPG and TD3 algorithms based on the action embedding method achieve similar or slightly better performance compared to the DQN-based algorithm. This is attributed to the large discrete action space, which makes it difficult for the DQN-based algorithm to explore adequately during training and to get the optimal policy after training. The DDPG algorithm reduces the long-term average cost of the greedy policy by $14.5$% from $826.2$ to $706.0$. TD3 is an extension of DDPG, which is designed to reduce overestimation bias and variance in the learning process by introducing delayed policy updates, target policy smoothing, and twin critic networks. These features can lead to overly conservative behavior in WNCS with high stochasticity, which potentially slows down the learning process or leads to underfitting where the system fails to adequately capture the beneficial actions amidst the noise. Thus, DDPG, as a more aggressive approach, is better than TD3.
@@
Fig. 4 and Fig. 5 showcase the comparison of three DRL algorithms and the three heuristic benchmark policies when the system scale is large. In these cases, the DQN-based algorithm is unable to solve the MDP, as the reduced discrete action space is still too large. Results show that the DDPG and TD3 algorithms based on the action embedding method effectively address the challenge of the rapidly growing action space, making it more suitable for large-scale WNCS problems. The DDPG algorithm reduces the long-term average cost of the greedy policy by $9.9$% from $602.0$ to $542.7$ when $N=10$ and $M=10$. As the communication resources become relatively scarce, the performance gain dramatically increases to $39.9$% (from $515.9$ to $309.6$), when $N=8$ and $M=6$, demonstrating the effectiveness of the proposed framework for more challenging scenarios.
Figure 5: The long-term average performance of DRL and benchmark algorithms over a system with $N=8$ and $M=6$. Values are means $\pm$ standard error of mean.
VI-C Practical Example
TABLE I: The Parameters of Pendulums
Pendulum #$i$ 1 2 3 4 5 6 7 8
Pole mass, $m_{p,i}$, [kg] 1 1.2 1.4 1.6 1.8 2 2.2 2.4
Pole length, $l_{p,i}$, [m] 1 1.1 1.2 1.3 1.4 1.5 1.6 1.7
We simulate a WNCS where 8 pendulums are spatially distributed and connected with a central controller via 6 frequency channels, as shown in Fig. 6(a). The parameters of these pendulums, including pole mass $m_{p,i}$ and pole length $l_{p,i}$ for each pendulum $i$, are shown in Table I. For pendulum $i$, the two-dimensional state vector $\mathbf{x}_{i,k}$, includes the pole angle state and the pole angular velocity state. The scaler control input $u_{i,k}$ is the applied torque. Each pendulum $i$ is linearized around the upright equilibrium using a small-angle approximation. The corresponding state transition matrix $\mathbf{A}_{i}$ and control input matrix $\mathbf{B}_{i}$ are given as [58]
$\mathbf{A}_{i}=\left[\begin{array}[]{cc}1&T_{s}\\ \frac{g}{l_{p,i}}T_{s}&1\end{array}\right],\mathbf{B}_{i}=\left[\begin{array}[]{c}T_{s}\\ \frac{T_{s}}{m_{p,i}l_{p,i}^{2}}\end{array}\right],$ (49)
where $T_{s}=0.01$ is the sampling time period; $g=9.81$ is the gravitational acceleration. The measurement matrix $\mathbf{C}_{i}$ is equal to the identity matrix. The covariance matrices are $\mathbf{Q}^{w}_{i}=\mathbf{Q}^{v}_{i}=0.001\mathbf{I}$. Each pendulum is $2$-step controllable, i.e., $v_{i}=2$.
The packet success probabilities of each link, i.e., $\xi^{+}_{m,i},\forall+\in\{s,c\}$, are simulated in the following. Let $\mathcal{F}_{c}\triangleq\{615\text{MHz},675\text{MHz},735\text{MHz},795\text{MHz},855\text{MHz},915\text{MHz}\}$ denote the available frequency bands. For links between each pendulum $i$ and the central controller, the free-space path loss model is $\bar{h}_{i}=A(\frac{3\times 10^{8}}{4\pi f_{i}d_{i}})^{d_{e}}$, where $A=4.11$ denotes the antenna gain; $f_{i}\in\mathcal{F}_{c}$ denotes the carrier frequency; $d_{i}$ denotes the distance from the pendulum to the central controller and is randomly drawn from $[180\text{m},200\text{m}]$; $d_{e}=2.2$ denotes the path loss exponent [59]. The average channel power gains for each link, $\mu^{s}_{m,i}$ and $\mu^{c}_{m,i}$, are randomly drawn from $[0.5,1.5]$. Given the transmission power $P_{\text{tx}}=23\text{dBm}$ and the receiving noise power $\sigma^{2}=-60\text{dBm}$, the signal-to-noise ratio (SNR) of received packets in each link is obtained from $\gamma_{m,i}^{+}=\frac{\bar{h}_{i}\mu^{+}_{m,i}P_{\text{tx}}}{\sigma^{2}}$.
Given the packet length $l=200$ (i.e., the number of symbols per packet), the number of data bits $b=400$ in the packet, and the SNR $\gamma_{m,i}^{+}$ of the packet, we have the approximated decoding error probability of a packet as [60]
$\varepsilon\left(\gamma_{m,i}^{+}\right)\approx\mathcal{Q}\left(\frac{\mathcal{C}\left(\gamma_{m,i}^{+}\right)-\frac{b}{l}}{\sqrt{\frac{\mathcal{V}\left(\gamma_{m,i}^{+}\right)}{l}}}\right),$ (50)
where $\mathcal{C}(\gamma_{m,i}^{+})=\log_{2}{(1+\gamma_{m,i}^{+})}$ and $\mathcal{V}(\gamma_{m,i}^{+})=(1-(1+\gamma_{m,i}^{+})^{-2})(\log_{2}{e})^{2}$ are the Shannon capacity and the channel dispersion, respectively, and $\mathcal{Q}(x)=(\frac{1}{\sqrt{2\pi}})\int_{x}^{\infty}{e^{-\frac{t^{2}}{2}}\text{d}t}$ is the Gaussian Q-function. Then, packet success probabilities of each link can be expressed as $\xi^{+}_{m,i}=1-\varepsilon\left(\gamma_{m,i}^{+}\right)$. The resulting WNCS satisfies the stability condition established in Theorem 1.
The DDPG-based scheduler is trained with $\mathbf{S}^{x}_{i}=\left[\begin{array}[]{cc}100&0\\ 0&1\end{array}\right]$, $\mathbf{S}^{u}_{i}=0.01$, and $\vartheta=0.95$. After training, the control performance is shown in Fig. 6(b). The pole angle of all pendulums is initialized at the down position. Results show that the pole angle of all pendulums can converge to zero, illustrating the effectiveness of the DDPG-based scheduler for stabilizing the pendulum system.
Figure 6: DDPG-based transmission scheduling for distributed control of the pendulum system: (a) Illustration of core components. (b) Testing results of the pole angle.
@@ VII Conclusions
We have investigated the transmission scheduling problem of the fully distributed WNCS. A sufficient stability condition of the WNCS in terms of both the control and communication system parameters has been derived. We have proposed advanced DRL algorithms to solve the scheduling problem, which perform much better than benchmark policies. For future work, we will develop a distributed DRL approach for enhancing the scalability of the scheduling algorithm. Furthermore, we will consider co-design problems of the estimator, the controller, and the scheduler for distributed WNCSs with time-varying channel conditions.